# Learning Motion Blur Robust Vision Transformers with Dynamic Early Exit for Real-Time UAV Tracking

### v1

@@ III-C Dynamic Early Exit Module (DEEM)
The primary goal of the DEEM is to allocate computational resources in a dynamic manner based on the complexity of each VIT block’s input, aiming to enhance inference efficiency while preserving performance. The DEEM operates by dynamically assessing the complexity of the input at each ViT block and making decisions regarding whether to continue processing or to exit early. This decision-making process is based on predefined criteria, specifically the cumulative exiting score exceeding a certain threshold. By exiting early for simpler inputs and reserving computational resources for more complex ones, the module optimizes the overall efficiency of the tracking system. While the primary objective of the DEEM is to improve inference efficiency, it does so without sacrificing tracking performance. The module is designed to ensure that early exits occur only when appropriate, thereby maintaining the accuracy and reliability of the tracking predictions. This balance between efficiency and performance is essential in real-time UAV tracking systems.
@@
The DEEM is accomplished by employing a linear layer in conjunction with a non-linear activation function, with a slice of all tokens representing both the target template and the search image as its input, as illustrated in Fig. 3. Its output represents the exiting score at the corresponding ViT block (layer). Specifically, take the $i$-th ($i>n_{enf}$) layer for example, the slice of all tokens ouput by the $(i-1)$-th ViT block is denoted by $\textbf{e}_{1}^{\textup{T}}\textbf{t}_{1:\mathcal{K}}^{i-1}(Z,X):=\textbf{b}^{i-1}\in\mathbb{R}^{\mathcal{K}}$, where $\textbf{e}_{1}^{\textup{T}}=[1,0,...,0]\in\mathbb{R}^{\mathcal{K}}$ is a standard unit vector in $\mathbb{R}^{\mathcal{K}}$, the linear layer is denoted by $\mathfrak{l}^{i}$, where $d$ is the token dimension. Formally the DEEM at layer $i$ is defined by $\mathfrak{e}^{i}=\mathfrak{s}(\mathfrak{l}^{i}(\textbf{b}^{i-1}))$, where $\mathfrak{e}^{i}\in[0,1]$ represents the exiting score of the $i$-th Transformer block, $\mathfrak{s}(x)=1/(1+e^{-x})$ denotes the sigmoid function. The forward pass exits if the cumulative exiting score $q^{i}=\sum_{k=n_{enf}+1}^{i}\lambda\mathfrak{e}^{k}$ exceeds $1-\epsilon$, i.e., $q^{i}\geqslant 1-\epsilon$, where $\epsilon$ is a small positive constant that allows exiting after examing one DEEM module. Note that, if $n_{enf}=0$, theoretically all $L$ blocks could be skipped simultaneously, resulting in no correlation computation between the template and search image. To prevent such an undesirable case, the first $n_{enf}>0$ layers are enforced without examination. This strategy is implemented to reduce computational burdens because these low-level blocks are expected in most cases, as low-level features provide foundational information upon which more complex features and representations can be built. Another extreme case is that, for any input, early exit never happens as by which the model is easier to reduce the training losses with more ViT blocks. To address this, we introduce a block sparsity loss, denoted as $\mathcal{L}_{spar}$, which penalizes smaller average exiting scores over all examined layers. This encourages, on average, an earlier exit during the forward pass to enhance efficiency. Let $L_{e}$ be the maximum layer of examined ViT blocks, it follows that
@@
$L_{e}=\underset{n_{enf}<n\leqslant L}{\textup{argmin}}\left\{\sum_{l=n_{enf}+1}^{n}\mathfrak{e}^{l}\geqslant 1-\varepsilon\right\}.$ (4)
@@
$\begin{split}\mathcal{L}_{spar}=||\frac{1}{L_{e}-n_{enf}}\sum^{L_{e}}_{l=n_{enf}+1}\mathfrak{e}^{l}-\tau||^{2},\end{split}$ (5)
@@
III-D Prediction Head and Training Objective
@@
TABLE I: Comparison of precision (Prec.), success rate (Succ.), and speed (FPS) between BDTrack and lightweight trackers on five UAV tracking benchmarks, namely DTB70, UAVDT, VisDrone2018, UAV123, and UAV123@10fps. Red, blue and green indicate the first, second and third place. It is worth noting that the percent symbol (%) has been omitted for Prec. and Succ. values.
DTB70 UAVDT VisDrone UAV123 UAV123@10fps Avg. Avg.FPS
Method Source Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. GPU CPU
KCF[7] TPAMI 15 46.8 28.0 57.1 29.0 68.5 41.3 52.3 33.1 40.6 26.5 53.1 31.6 - 622.5
BACF[44] ICCV 17 58.1 39.8 68.6 43.2 77.4 56.7 66.0 45.9 57.2 41.3 65.5 45.4 - 54.2
fDSST[45] TPAMI 17 53.4 35.7 66.6 38.3 69.8 51.0 58.3 40.5 51.6 37.9 60.0 40.7 - 193.4
ECO_HC[46] CVPR 17 63.5 44.8 69.4 41.6 80.8 58.1 71.0 49.6 64.0 46.8 69.7 48.2 - 83.5
MCCT_H[47] CVPR 18 60.4 40.5 66.8 40.2 80.3 56.7 65.9 45.7 59.6 43.4 66.6 45.3 - 63.4
STRCF[48] CVPR 18 64.9 43.7 62.9 41.1 77.8 56.7 68.1 48.1 62.7 45.7 67.3 47.1 - 28.4
ARCF[9] ICCV 19 69.4 47.2 72.0 45.8 79.7 58.4 67.1 46.8 66.6 47.3 71.0 49.1 - 34.2
AutoTrack[11] CVPR 20 71.6 47.8 71.8 45.0 78.8 57.3 68.9 47.2 67.1 47.7 71.6 49.0 - 57.8
DCF-based RACF[6] PR 22 72.6 50.5 77.3 49.4 83.4 60.0 70.2 47.7 69.4 48.6 74.6 51.2 - 35.6
HiFT[1] ICCV 21 80.2 59.4 65.2 47.5 71.9 52.6 78.7 59.0 74.9 57.0 74.2 55.1 160.3 -
SiamAPN++[49] IROS 21 78.9 59.4 76.9 55.6 73.5 53.2 76.8 58.2 76.4 58.1 76.5 56.9 167.5 -
LightTrack[50] CVPR 21 76.4 59.1 80.4 61.2 74.8 56.8 78.3 62.7 75.1 59.9 77.0 59.9 121.4 -
HCAT[58] ECCV 22 83.5 63.7 74.2 54.3 75.4 57.7 83.0 67.8 81.9 64.4 79.6 61.6 144.4 -
TCTrack[5] CVPR 22 81.2 62.2 72.5 53.0 79.9 59.4 80.0 60.5 78.0 59.9 78.3 59.0 139.6 -
UDAT[55] CVPR 22 80.6 61.8 80.1 59.2 81.6 61.9 76.1 59.0 77.8 58.5 79.2 60.1 33.7 -
SGDViT[56] ICRA 23 78.5 60.4 65.7 48.0 72.1 52.1 75.4 57.5 86.3 66.1 75.6 56.8 110.5 -
ABDNet[33] RAL 23 76.8 59.6 75.5 55.3 75.0 57.2 79.3 60.7 77.3 59.1 76.7 59.1 130.2 -
CNN-based DRCI[57] ICME 23 81.4 61.8 84.0 59.0 83.4 60.0 76.7 59.7 73.6 55.2 79.8 59.1 281.3 62.7
Aba-ViTrack[18] ICCV 23 85.9 66.4 83.4 59.9 86.1 65.3 86.4 66.4 85.0 65.5 85.3 64.7 181.5 50.3
HiT[51] ICCV 23 75.1 59.2 62.3 47.1 74.8 58.7 80.6 63.8 80.9 64.3 74.7 58.6 237.7 57.3
LiteTrack[52] arXiv’23 82.5 63.9 81.6 59.3 79.7 61.4 84.2 65.9 83.1 65.0 82.2 63.1 141.6 -
SMAT[53] WACV 24 81.9 63.8 80.8 58.7 82.5 63.4 81.8 64.6 80.4 63.5 81.5 62.8 124.2 -
LightFC[54] KBS 24 82.8 64.0 83.4 60.6 82.7 62.8 84.2 65.5 81.3 63.7 82.9 63.4 146.8 -
BDTrack-ViT 82.8 64.2 78.9 57.3 83.9 63.6 83.5 65.5 84.1 66.1 82.6 63.3 235.1 56.9
BDTrack-T2T 80.7 63.1 77.9 56.2 81.4 62.2 80.4 63.4 82.9 65.1 80.7 62.0 224.4 55.8
VIT-based BDTrack-DeiT Ours 83.5 64.1 84.1 61.0 85.2 64.3 84.8 66.7 83.5 65.9 84.2 64.4 287.2 63.9
IV Experiments
@@
This section provides detailed and comprehensive evaluation results of our approach across five well-known UAV tracking benchmarks: namely, UAV123 [61], UAV123@10fps [61], VisDrone2018 [62], UAVDT [63], and DTB70 [64]. A PC equipped with an NVIDIA TitanX GPU, 16GB RAM, and an i9-10850K CPU (3.6GHz) was used for all evaluation experiments. In order to provide a comprehensive evaluation, we evaluate our method against 23 lightweight SOTA trackers, which are categorized into three groups as in Aba-ViTrack, i.e., DCF-based, CNN-based, and ViT-based methods (see Table I), alongside 12 SOTA deep trackers tailored for generic visual tracking (refer to Table II).
IV-A Implementation Details
@@
We employed three efficient ViT models, i.e., ViT-tiny [65], DeiT-tiny [66], and T2T-tiny [67], as the backbones to develop three distinct trackers, named BDTrack-ViT, BDTrack-DeiT, and BDTrack-T2T, respectively. The sizes of the search region and template are set to $256^{2}$ and $128^{2}$. The head architecture, training data, hyperparameters, and training pipeline adhere to the specifications outlined in Aba-ViTrack [18]. And we apply Hanning window penalties during inference to integrate positional priors in tracking, following established practices [68].
IV-B Comparison with Lightweight Trackers
@@
In this part, we thoroughly evaluate the performance of our BDTrack by comparing it with 23 cutting-edge lightweight trackers, including KCF [7], BACF [44], fDSST [45], ECO_HC [46], MCCT_H [47], STRCF [48], ARCF [9], AutoTrack [11], RACF [6], HiFT [1], SiamAPN++ [49], LightTrack [50], HCAT [58], TCTrack [5], UDAT [55], SGDViT [56], ABDNet [33], DRCI [57], Aba-ViTrack [18], HiT [51], LiteTrack [52], SMAT [53], and LightFC [54], on five UAV tracking benchmarks. The evaluation results are shown in Table I. As can be seen, our BDTrack-DeiT outperforms other trackers except Aba-ViTrack on all benchmarks, regarding average (Avg.) precision (Prec.) and success rate (Succ.). RACF [6] secures the highest Prec.and Succ. in average among DCF-based trackers, with values of 74.6% and 51.2%, respectively. Among CNN-based trackers, DRCI [57] stands out with the highest average precision of 79.8%, while HCAT [58] achieves the highest success rate of 61.6%. However, all ViT-based trackers exhibit Avg. Prec. and Succ. exceeding 80.0% and 62.0%, respectively, except for HiT, obviously beating the CNN-based methods and substantially outperforming the DCF-based ones. In terms of GPU speed, BDTrack-DeiT boasts the highest speed at 287.2 FPS. DRCI and HiT follow closely, achieving the second and third places with speeds of 281.3 FPS and 237.7 FPS, respectively. Despite its comparable GPU speed to BDTrack-DeiT, DRCI exhibit significantly lower Avg. Prec. and Succ. compared to BDTrack-DeiT. As for CPU speed, all our trackers exhibit real-time performance on a single CPU [^1] Note that the real-time performance discussed in this paper can be generalized only to platforms similar to or more advanced than ours.. However, it is noteworthy that all methods achieving speeds above 80 FPS belong to the DCF-based trackers, suggesting that the most efficient UAV trackers are still DCF-based. Although Aba-ViTrack achieves the best average performance with 85.3% Avg. Prec. and 64.7% Avg. Succ., our BDTrack-DeiT secures the second position with only slight gaps of 1.1% and 0.3%, respectively. Notably, BDTrack-DeiT surpasses Aba-ViTrack in Prec. on UAVDT and in Succ. on both UAVDT and UAV123. Furthermore, all our trackers exhibit higher speeds than Aba-ViTrack. Remarkably, BDTrack-DeiT boasts over 1.5 times the GPU speed and 1.2 times the CPU speed of Aba-ViTrack, striking a better balance between tracking precision and efficiency. These results highlight the advantages of our method and justify its SOTA performance for UAV tracking.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: The precision plots and success plots of attribute-based comparison are presented for the attribute subsets of UAVDT. Note that BDTrack-DeiT* denotes BDTrack-DeiT without the application of the proposed components.
IV-C Attribute-Based Evaluation
To further validate the superiority of the proposed tracker over existing UAV trackers, we assess the performance of BDTrack-DeiT by comparing it with 14 SOTA UAV trackers on three subsets of the UAVDT dataset, including ‘Motion blur’, ‘Camera motion’, and ‘Scale variations’. Note that we also evaluate BDTrack-DeiT without applying the components (i.e., MBRV and DEEM), denoted as BDTrack-DeiT* for reference. The precision and success plots are illustrated in Fig. 4. As observed, BDTrack-DeiT significantly outperforms the second tracker on ‘Motion blur’, ‘Camera motion’ and ‘Scale variations’, with gains of 1.2%/0.4%, 2.1%/0.7%, and 3.5%/3.0% in precision/success rate, respectively. Significantly, across all these three attributes, the incorporation of the proposed components yields a noteworthy enhancement over BDTrack-DeiT*, with improvements of 7.0%/5.0%, 7.9%/5.5%, and 11.7%/9.1% in terms of precision and success rate. Note that ”Camera motion” and ”Scale variations” are both factors that can contribute to motion blur in visual tracking systems. When a camera moves rapidly or changes direction suddenly, it can cause motion blur in the captured images. While scale variations themselves do not directly cause motion blur, they can complicate the tracking process. If an object moves quickly closer or farther away and the camera or algorithm doesn’t adjust well, the object can appear blurred in the images. This happens because the object’s speed relative to the camera’s frame rate and exposure time changes, which can result in motion blur. Therefore, this attribute-based evaluation validates, directly or indirectly, the superiority and effectiveness of the proposed method in addressing the challenges posed by motion blur.
Figure 5: Qualitative evaluation on 4 video sequences from DTB70, UAVDT, UAV123@10fps, and VisDrone2018 (i.e. RcCar7, S0601, uav2, and uav0000208_00000_s), respectively.
IV-D Qualitative evaluation
Qualitative tracking results of BDTrack-DeiT and seven SOTA trackers, i.e., Aba-ViTrack [18], HiFT [1], TCTrack [5], ABDNet [33], RACF [6], UDAT [55], and AutoTrack [11], are presented in Fig. 5. Four video sequences from four different benchmarks are selected for demonstration: RcCar7 from DTB70, S0601 from UAVDT, uav2 from UAV123@10fps, and uav0000208_00000_s from VisDrone2018. For better visualization, the low-resolution targets within the video frames are selectively zoomed, cropped, and placed in the upper right corner of their corresponding frames for display. Notably, only our tracker successfully tracks the targets in all challenging examples, featuring motion blur (i.e, RcCar7 and uav2), occlusion (i.e, S0601 and uav0000208_00000_), low resolution (i.e, uav2), and scale variations (i.e, in all sequences). Specifically, BDTrack-DeiT is the only tracker to successfully track the target in uav0000208_00000_s. In S0601, while Aba-ViTrack, RACF, AutoTrack, and BDTrack-DeiT all successfully track the target, BDTrack-DeiT demonstrates higher accuracy. In the remaining sequences, BDTrack-DeiT consistently proves to be the most accurate tracker. Our method not only exhibits superior performance but also delivers more visually pleasing results in these scenarios, further validating the effectiveness of the proposed method for UAV tracking.
IV-E Comparison with Deep Trackers

### v2

@@ 3.3 Dynamic Early Exit Module (DEEM)
The primary goal of the DEEM is to allocate computational resources in a dynamic manner based on the complexity of each VIT block’s input, aiming to enhance inference efficiency while preserving performance. The DEEM operates by dynamically assessing the complexity of the input at each ViT block and making decisions regarding whether to continue processing or to exit early. This decision-making process is based on predefined criteria, specifically the cumulative exiting score exceeding a certain threshold. By exiting early for simpler inputs and reserving computational resources for more complex ones, the module optimizes the overall efficiency of the tracking system. While the primary objective of the DEEM is to improve inference efficiency, it achieves this with minimal impact on tracking performance. The module is designed to ensure that early exits occur only when appropriate, thereby maintaining the accuracy and reliability of the tracking predictions. This balance between efficiency and performance is essential in real-time UAV tracking systems.
@@
The DEEM is accomplished by employing a linear layer in conjunction with a non-linear activation function, with a slice of all tokens representing both the target template and the search image as its input, as illustrated in Fig. 3. Its output represents the exiting score at the corresponding ViT block (layer). Specifically, take the $i$-th ($i>n_{enf}$) layer for example, the slice of all tokens ouput by the (i1)(i-1)-th ViT block is denoted by $\textbf{e}_{1}^{\textup{T}}\textbf{t}_{1:\mathcal{K}}^{i-1}(Z,X):=\textbf{b}^{i-1}\in\mathbb{R}^{\mathcal{K}}$, where $\textbf{e}_{1}^{\textup{T}}=[1,0,...,0]\in\mathbb{R}^{\mathcal{K}}$ is a standard unit vector in $\mathbb{R}^{\mathcal{K}}$, the linear layer is denoted by $\mathfrak{l}^{i}$, where $d$ is the token dimension. Formally, the DEEM at layer $i$ is defined by $\mathfrak{e}^{i}=\sigma(\mathfrak{l}^{i}(\textbf{b}^{i-1}))$, where $\mathfrak{e}^{i}\in[0,1]$ represents the exiting score of the $i$-th Transformer block, $\sigma(x)=1/(1+e^{-x})$ denotes the sigmoid function. The forward pass exits if the cumulative exiting score $q^{i}=\sum_{k=n_{enf}+1}^{i}\lambda\mathfrak{e}^{k}$ exceeds $1-\epsilon$, i.e., $q^{i}\geqslant 1-\epsilon$, where $\epsilon$ is a small positive constant that enables early exiting after examining a single DEEM module. In our implementation, we set $\epsilon=0.05$, which allows an early exit when the cumulative confidence exceeds 0.95. Note that, if $n_{enf}=0$, theoretically all $L$ blocks could be skipped simultaneously, resulting in no correlation computation between the template and search image. To prevent such an undesirable case, the first $n_{enf}>0$ layers are enforced without examination. This strategy is implemented to reduce computational burdens because these low-level blocks are expected in most cases, as low-level features provide foundational information upon which more complex features and representations can be built. Another extreme case is that, for any input, early exit never happens as by which the model is easier to reduce the training losses with more ViT blocks. To address this, we introduce a block sparsity loss, denoted as $\mathcal{L}_{spar}$, which penalizes smaller average exiting scores over all examined layers. This encourages, on average, an earlier exit during the forward pass to enhance efficiency. Let $L_{e}$ be the maximum layer of examined ViT blocks, it follows that
@@
$L_{e}=\underset{n_{enf}<n\leqslant L}{\textup{argmin}}\left\{\sum_{l=n_{enf}+1}^{n}\mathfrak{e}^{l}\geqslant 1-\epsilon\right\}.$ (4)
@@
[…]
@@
3.4 Prediction Head and Training Objective
@@
Table 1: Comparison of precision (Prec.) and success rate (Succ.) between BDTrack and lightweight trackers on four UAV tracking benchmarks, namely UAVDT [58], VisDrone2018 [59], UAV123 [60], and UAV123@10fps [60]. Red, blue and green indicate the first, second and third place. It is worth noting that the percent symbol (%) has been omitted for Prec. and Succ. values.
UAVDT VisDrone2018 UAV123 UAV123@10fps Avg.
Method Source Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ.
KCF [22] TPAMI 15 57.1 29.0 68.5 41.3 52.3 33.1 40.6 26.5 54.6 32.5
fDSST [61] TPAMI 17 66.6 38.3 69.8 51.0 58.3 40.5 51.6 37.9 61.6 41.9
BACF [62] ICCV 17 68.6 43.2 77.4 56.7 66.0 45.9 57.2 41.3 67.3 46.8
ECO_HC [63] CVPR 17 69.4 41.6 80.8 58.1 71.0 49.6 64.0 46.8 71.3 49.0
MCCT_H [64] CVPR 18 66.8 40.2 80.3 56.7 65.9 45.7 59.6 43.4 68.2 46.5
ARCF [23] ICCV 19 72.0 45.8 79.7 58.4 67.1 46.8 66.6 47.3 71.4 49.6
AutoTrack [24] CVPR 20 71.8 45.0 78.8 57.3 68.9 47.2 67.1 47.7 71.7 49.3
DCF-based RACF [21] PR 20 77.3 49.4 83.4 60.0 70.2 47.7 69.4 48.6 75.1 51.4
HiFT [7] ICCV 21 65.2 47.5 71.9 52.6 78.7 59.0 74.9 57.0 72.7 54.0
SiamAPN++ [17] IROS 21 76.9 55.6 73.5 53.2 76.8 58.2 76.4 58.1 75.9 56.3
LightTrack [65] CVPR 21 80.4 61.2 74.8 56.8 78.3 62.7 75.1 59.9 77.2 60.2
TCTrack [8] CVPR 22 69.1 50.4 77.6 57.7 77.3 60.4 75.1 58.8 74.7 56.8
TCTrack++ [66] TPAMI 23 72.5 53.2 80.8 60.3 74.4 58.8 78.2 60.1 76.5 58.1
SGDViT [67] ICRA 23 65.7 48.0 72.1 52.1 75.4 57.5 86.3 66.1 74.9 55.9
ABDNet [44] RAL 23 75.5 55.3 75.0 57.2 79.3 60.7 77.3 59.1 76.8 58.1
DRCI [68] ICME 23 83.0 59.0 83.4 60.0 76.7 59.7 73.6 55.2 79.2 58.5
CNN-based PRL-Track [69] IROS 24 73.1 53.5 72.6 53.8 79.1 59.3 74.1 58.6 74.7 56.3
Aba-ViTrack [9] ICCV 23 83.4 59.9 86.1 65.3 86.4 66.4 85.0 65.5 85.2 64.3
LiteTrack [70] ICRA 24 81.6 59.3 79.7 61.4 84.2 65.9 83.1 65.0 82.2 62.9
SMAT [71] WACV 24 80.8 58.7 82.5 63.4 81.8 64.6 80.4 63.5 81.4 62.6
LightFC [72] KBS 24 83.4 60.6 82.7 62.8 84.2 65.5 81.3 63.7 82.9 63.1
AVTrack-DeiT [73] ICML 24 82.1 58.7 86.0 65.3 84.8 66.8 83.2 65.8 84.0 64.1
BDTrack-ViT
78.9 57.3 83.9 63.6 83.5 65.5 84.1 66.1 82.6 63.1
ViT-based
BDTrack-DeiT
Ours 84.1 61.0 85.2 64.3 84.8 66.7 83.5 65.9 84.4 64.5
4 Experiments
@@
This section provides detailed and comprehensive evaluation results of our approach across four well-known UAV tracking benchmarks: namely, UAV123 [60], UAV123@10fps [60], VisDrone2018 [59], and UAVDT [58]. A PC equipped with a NVIDIA TitanX GPU, 16GB RAM, and an i9-10850K CPU (3.6GHz) was used for all evaluation experiments. In order to provide a comprehensive evaluation, we evaluate our method against 22 lightweight SOTA trackers, which are categorized into three groups as in Aba-ViTrack [9], i.e., DCF-based, CNN-based, and ViT-based methods (see Table 1), alongside 14 SOTA deep trackers tailored for generic visual tracking (refer to Table 2). Note that in our study, “light trackers" and “efficient trackers" are used interchangeably to refer to trackers that prioritize low computational complexity and reduced memory usage, typically designed for environments with limited resources or real-time applications.
4.1 Implementation Details
@@
We employed three efficient ViT models, i.e., ViT-tiny [74] and DeiT-tiny [75], as the backbones to develop three distinct trackers, named BDTrack-ViT, BDTrack-DeiT, and BDTrack-T2T, respectively. The sizes of the search region and template are set to $256^{2}$ and $128^{2}$. The head architecture, training data, hyperparameters, and training pipeline adhere to the specifications outlined in Aba-ViTrack [9]. And we apply Hanning window penalties during inference to integrate positional priors in tracking, following established practices [76].
4.2 Comparison with Lightweight Trackers
@@
In this section, we comprehensively evaluate BDTrack’s performance by comparing it with 22 SOTA lightweight trackers across four UAV tracking benchmarks. The evaluation results are shown in Table 1. As can be seen, our BDTrack-DeiT outperforms other trackers except Aba-ViTrack on all benchmarks, regarding average (Avg.) precision (Prec.) and success rate (Succ.). RACF [21] secures the highest Prec. and Succ. in average among DCF-based trackers, with values of 75.1% and 51.4%, respectively. Among CNN-based trackers, DRCI [68] stands out with the highest average Prec. of 79.2% and Succ. of 58.5%,. However, all ViT-based trackers exhibit Avg. Prec. and Succ. exceeding 81.0% and 62.0%, respectively, obviously beating the CNN-based methods and substantially outperforming the DCF-based ones. Although Aba-ViTrack achieves the highest Avg. Prec. (85.2%), our BDTrack-DeiT ranks second with only a slight gap of 0.8%, while maintaining the best Avg. Succ. of 64.5%. Notably, BDTrack-DeiT outperforms Aba-ViTrack on the UAVDT, with a success rate improvement of over 1.0%. These results highlight the advantages of our method and justify its SOTA performance for UAV tracking.
4.3 Comparison with Deep Trackers